OpenStudy (anonymous):

solve this DE: y'+y=e^{3x}

OpenStudy (anonymous):

\[e^{\int{}{}p(x)dx}\]

OpenStudy (anonymous):

\[e^{\int{}{}dx}=e^{x+c}\]

OpenStudy (anonymous):

o wait i know what i didn' do

OpenStudy (unklerhaukus):

you dont need a contant of integration 'c' in your integrating factor

OpenStudy (unklerhaukus):

\[\frac{\text d}{\text d x}(ye^x)=e^x\cdot e^{3x}\]

OpenStudy (anonymous):

yep then i forgot to integrate the right side for some reason. i think it's because the question before was a homogeneous equation

OpenStudy (unklerhaukus):

you can finish from here?

OpenStudy (anonymous):

yes